OpenStudy (anonymous):

- 24/55 divided by 39/40

OpenStudy (asnaseer):

use the fact that:\[\frac{a}{b}\div\frac{c}{d}=\frac{a}{b}\times\frac{d}{c}\]

OpenStudy (anonymous):

do i cross multiply

OpenStudy (asnaseer):

rewrite the division as a multiplication using the rule above and then cancel common factors between the numerators and the denominators.

OpenStudy (anonymous):

how do i do that

OpenStudy (asnaseer):

I suggest you go through this site to learn about how to manipulate fractions: http://www.aaamath.com/fra.html#topic23

OpenStudy (anonymous):

well i only ask how to set this up i never ask you to give me the answer.

OpenStudy (asnaseer):

I have given you the information above. That site will help you understand how to do these types of questions as it also hows you lots of examples.

OpenStudy (asnaseer):

*shows

OpenStudy (anonymous):

ok thank you!

OpenStudy (asnaseer):

in your case you have:\[\frac{24}{55}\div\frac{39}{40}=\frac{24}{55}\times\frac{40}{39}\]

OpenStudy (anonymous):

see the thing was i already had it up to this point I just wasnt sure what you meant for me to do next do i find the GCF of both? because if so i know how to do at least that.

OpenStudy (asnaseer):

no - you just need to cancel out common factors between numerators and denominators. e.g. the following fraction can be simplified as follows:\[\frac{12}{28}=\frac{3}{7}\]because I divided both the numerator and the denominator by their common factor (4 in this case)

OpenStudy (asnaseer):

remember that:\[\frac{24}{55}\times\frac{40}{39}=\frac{24*40}{55*39}\]so, for example, I can see that the 40 in the numerator and the 55 in the denominator can both be divided by 5 to get:\[\frac{24*40}{55*39}=\frac{24*8}{11*39}\]keep dividing the numerator and denominators by any common factors you can find.

OpenStudy (anonymous):

ok thank! this website you told me that would help is it free?

OpenStudy (asnaseer):

yes

OpenStudy (asnaseer):

This might be better: http://www.mathsisfun.com/fractions_division.html

OpenStudy (anonymous):

ok thanks!

OpenStudy (asnaseer):

yw

OpenStudy (anonymous):

one more thing when i turn the fracton into a mulitiplication what happens to the negative sign

OpenStudy (asnaseer):

oh - I didn't realise you had a negative sign. the negative sign just gets placed in front of the overal fraction as follows:\[-\frac{24}{55}\times\frac{40}{39}=-\frac{24*40}{55*39}\]

OpenStudy (anonymous):

would it be - 64/143

OpenStudy (asnaseer):

perfect!

OpenStudy (anonymous):

this is very had did you see how long it took me to find this out lol

OpenStudy (asnaseer):

don't worry - we all had to start somewhere - with practice you will build confidence in these types of questions and I'm sure you'll be doing them in your sleep very soon! :)

OpenStudy (anonymous):

trust me im far from that lol

OpenStudy (asnaseer):

:D

OpenStudy (asnaseer):

well - if you need help you only need to ask :)